The distribution of iodine between carbon tetrachloride and water is constant. The value of K is cccia / cH2o = 85. How many ml of carbon tetrachloride would be required to extract 959% iodine from 100 ml of aqueous solution in a single phase?



1
Expert's answer
2021-04-05T07:45:13-0400

K = Cccl4 / CH20 = 85

If we have 1mmol/L Solution of I2 in water, 0.959 mol must be extracted and 0.041 mmol left in water

C = n/V

C(Iodine in water) = 0.041/0.1 = 0.41mol/L

85 = X/0.041, X = 0.41× 85 = 34.85 mmol/L

VCcl4 = 0.959/34.85 = 0.0275 L

= 27.5 ml